Given: A particle is moving along a straight line such that its velocity is defined as v = (-4s2) m/s, where s is

in meters.

Find: The velocity and acceleration as functions of time if s = 2 m when t = 0.


1
Expert's answer
2020-09-15T09:43:25-0400

"V= \frac {ds}{dt}"


"V=-4s^2 m\/s"


"\frac {ds}{dt}=-4s^2"


"S= \frac {1}{4t+ \frac{1}{2}}" "=\frac {2} {8t+1}m"


"v=\frac{ds}{dt}= \frac {d}{dt}" "=(\frac {2}{8t+1})" "=" "=\frac {-16}{(8t+1)^2}"


"velocity= \frac {-16}{(8t+1)^2}"


"a=\frac{dv}{dt}" = "\frac{16(2)(8t+1)(8)}{(8t+1)^4}"


"a= \frac {256}{(8t+1)^3} m/s^2"